OpenStudy (anonymous):

solve the equation by clearing fractions. -2 + 3/4 x=5x -1/2

OpenStudy (anonymous):

-3/2=17/4 x -6=17x -6/17=x

OpenStudy (anonymous):

where did the 17 come from?

OpenStudy (anonymous):

5x=20/4 x 20/4 x -3/4 x=17/4 x